# The PLATO Mission

### v1

@@ 13.2 PLATO Cameras
Thermally, each camera is individually controlled by a Thermal Control System (TCS) that sets its temperature thanks to three heaters placed on the tube of the TOU. Using a Proportional-Integral (PI) controller algorithm, the temperature of each camera can be held very stable, up to mK over 14 hours using 600 s sliding window averages. Thanks to the athermal design of the camera and the AlBeMet tube properties, this temperature stability around the heaters is very well propagated to all the camera sub-systems and especially the CCDs. The other way around, the heat dissipated by the CCDs is conductively transferred directly to the baffle that can then dissipate it into cold space. This TCS is also used to fine tune the cameras focus in space and make sure that each individual camera is being used in its best possible focus position [247, 82]. By changing the temperature of the camera, the characteristics of the optical system are changing leading to a slight shift of the focus. This thermal refocusing is used on-ground and then during commissioning to calibrate each camera individually to its best focus temperature. Testing under representative thermal vacuum conditions on the EM1, the first complete, flight-like camera, has shown that the best focus will be around C. The precise value for each camera will be determined during the commissioning phase once in space.
The optical design of each camera [200] consists of a six lens system with a central Calcium Fluoride lens [41] close to the stop, and with a first aspheric lens protected, for thermal and radiation hardness purposes, by a Suprasil window. All the optical materials have been the subject of specific radhard studies [89] and the whole design is a fine tuning balance between radiation, thermal and mass budgets [201].
An anti-reflection coating is placed on every optical surface to maximise the optical transmission in the spectral range [500-1000] nm. Each 12 cm aperture camera has a FoV of 1037 square degrees as shown in [248]. This wide FoV is achieved thanks to a circular optical FoV of the TOU of around 18.88° radius that illuminates the 4 full-frame CCDs of 81.18 81.18 mm² each on the Focal Plane Assemblies (FPA). As mentioned above, each FPA is electrically connected to a Front End Electronics (either "normal" or "fast" ones), completing the PLATO Camera. One of the FPA models used for qualification is shown in Fig. 28, with 4 full-frame (i.e. "normal" CCDs). More details on the FPA can be found in Moreno et al [224]. Table 8 summarises the the main parameters of the individual cameras.
@@ 13.3 PLATO Data Processing System
The PLATO payload on-board data processing system is made up of an Instrument Control Unit (ICU), two Main Electronic Units (MEU) and a Fast Electronic Unit (FEU). Each MEU contains 6 normal Data Processing Units (N-DPUs) for processing data from the normal cameras. Each N-DPU is dealing with two normal cameras in terms of managing the N-FEE and acquiring data. The nominal processing cadence for the N-DPUs is 25 sec. There are two F-DPUs gathered in the FEU. Each F-DPU is responsible for processing the data of one fast camera. The processing cadence for F-DPUs is 2.5 sec. The F-DPUs main purpose is providing attitude data for the Fine Guidance System (FGS) directly to the Service Module (SVM) AOCS and to the ICU. Furthermore, it manages the F-FEE and handles scientific data of bright stars. There are two ICU channels which work in cold redundancy. The ICU is responsible for the management of the payload, the communication with the SVM and the compression of scientific data before transmitting them as telemetry to the SVM. Data is routed through a SpaceWire network from FEU and MEU to the ICU, and then from the ICU to the SVM. For the FGS data, dedicated SpaceWire links between FEU and spacecraft AOCS are used. Figure 29 gives an overview of the PLATO data processing system architecture and of the data flow rates while Fig. 30 shows the engineering models during the tests at DLR. It focuses on the sharing of the main functions and the data flows. It is a simplified view of the hardware architecture.
@@ 14 Synergies with other Missions
Exoplanet missions form a significant part of ESA and NASA’s ongoing and future space science programs ([286] for an overview). PLATO will start its science operations in 2027. By this time, the extended operations periods for the now ongoing missions CHEOPS and TESS will have ended, although further prolongations could extend their lifetimes until and beyond PLATO launch. Prolongation of James-Webb-Space-Telescope (JWST) operations beyond 2026 will allow it to observe PLATO targets.Two years after PLATO, the ARIEL mission will launch to spectroscopically observe transiting exoplanets. Here we outline the key aspects of synergies with PLATO for these and other missions:
@@
TESS: A discussion of the planet yield of PLATO versus TESS [259] can be found in Section 2. The majority of planets detected by TESS have orbital periods 10 days [135], while PLATO planets focuses on orbital periods 27 days, including a significant fraction on long-periods 100 days. Most of the planet detections around solar-like stars from PLATO will have , outnumbering the numbers expected from TESS for this type of stars. There is a clear complementarity of TESS and PLATO.
@@
CHEOPS: The main goal is to search for transits on bright stars already known to host planets to determine precise radii [39]. A major benefit for PLATO from the CHEOPS mission results from improved understanding and modelling activities in the community on the interpretation of precise planet radii, masses and ages in terms of, e.g, planet formation, evolution and interior atmosphere interactions. While many of the related science questions can only be fully answered when a larger sample of accurately characterised planets from PLATO becomes available, the ongoing activities within CHEOPS pave the way for PLATO data to come. Although likely only possible for very few targets, a cross-calibration of CHEOPS with PLATO will be beneficial. Furthermore, where applicable preparatory observations of PLATO targets by CHEOPS can be beneficial to maximise science return on these objects (e.g. for TTV studies).
ARIEL: will provide spectra of atmospheres of a large number of transiting planets [294]. These planets will mostly be hot and warm Jupiters, with some Neptunes and super-earths around late-type stars. We expect that (ultra-)hot Neptunes and larger target objects on orbits 10 days suitable for ARIEL will already be known via TESS (see Section 2). Based on this, about 20-40 targets are expected in potential PLATO LOPs. Additional targets on smaller, warm planets around bright stars newly detected by PLATO will form additional exciting targets for ARIEL. PLATO will provide high-accuracy radii, prolonged TTV data and well-known ephemeris for these planets within its target fields. Information from accurate N-CAM white-light phase curves and two-colour information from PLATO´s F-CAMs for hot and nearby targets will help to select especially interesting targets for ARIEL (see Section 3.3). Input of PLATO on well-known radii, masses and stellar parameters, including the age, will be beneficial for ARIEL for targets in the PLATO fields. For example, PLATO ages will help breaking the degeneracies for warm Jupiter planets. How planet properties correlate with age (e.g. due to contraction, atmosphere loss, tidal interaction, etc.) can be studied once ages are available in sufficient numbers and accuracy from PLATO. In addition, the pre-characterisation of stellar activity by PLATO (rotational modulation, phase of activity cycle, frequency of flares and CMEs, granulation) will be helpful. Such data will be especially useful for simultaneously observed targets.
@@
GAIA: Through its astrometric survey of the Milky Way, Gaia [117] will discover a large population of massive long period planets, primarily astrometrically. Given that orbital inclinations are known for astrometrically detected planets, PLATO might be able to identify transits on some of them, where even single transit detections (mono-transits) would be very useful and could become the first transit observations of long-period planets. Furthermore, for the vetting of transiting planet candidates, GAIA will help in the identification of background ecplising binaries as has been demonstrated for TESS candidates by [243]. A significant harvest of new long period ( 5 years), massive exoplanets is expected in the upcoming Gaia data releases (e.g. Gaia DR4).The combination of long-period Gaia planet detections with the smaller and intermediate orbital distance exoplanets discovered by PLATO, will enable the discovery of exoplanets systems with architectures analogous to our own, namely exoplanet systems with shorter period ‘rocky planets’ and longer period gas giants orbiting solar types stars.
Euclid and Roman Space telescope: The Nancy Grace Roman Space telescope [285] is expected to detect exoplanets from microlensing, from transits and from coronographic imaging, covering a wide range of planet system parameters. ESA’s Euclid mission also has some capability to detect extrasolar planets, and studies demonstrating the joint detection capabilities of the two missions have been performed (e.g. Bachelet et al 19, Kerins et al 173). Correlations of planet population statistics derived from PLATO with Euclid and the Roman Telescope results will provide valuable inputs into further constraining planet formation.
@@ 15 Appendix A
where describes the effect of limb darkening. The current knowledge of limb darkening for low magnetic field, quiet, solar-type stars is known sufficiently well that the contribution to the uncertainty in the radius ratio is negligible (c.f. Section 4.4. of Maxted 208 as well as Ludwig et al 197). However, numerical simulations show that about over 800 Gauss surface magnetic field the theory predicts higher limb darkening than that was observed. In the same way, some simulations result in limb brightening effects difficult to understand at this point (see Ludwig et al 197). So, the very active stars can have different properties and more work is needed to calibrate properly their limb darkening behaviour (c.f. Csizmadia et al 91, Espinoza and Jordán 110, Agol et al 6, Patel and Espinoza 246). The results are not changing if we consider different limb darkening laws; therefore, for sake of simplicity, we used linear limb darkening with a constant linear limb darkening coefficent which is a solar-like value (). Then we have by differentiation that

### v2

@@ 13.2 PLATO Cameras
Thermally, each camera is individually controlled by a Thermal Control System (TCS) that sets its temperature thanks to three heaters placed on the tube of the TOU. Using a Proportional-Integral (PI) controller algorithm, the temperature of each camera can be held very stable, up to mK over 14 hours using 600 s sliding window averages. Thanks to the athermal design of the camera and the AlBeMet tube properties, this temperature stability around the heaters is very well propagated to all the camera sub-systems and especially the CCDs. The heat dissipated by the CCDs is conductively transferred directly to the baffle that can then dissipate it into cold space. This TCS is also used to fine tune the focus of each camera in space and ensure that each camera is operated in its best possible focus position [270, 87]. By changing the temperature of the camera, the characteristics of the optical system are changing leading to a slight shift of the focus. This thermal refocusing is used on-ground and then during commissioning to calibrate each camera individually to its best focus temperature. Testing under representative thermal vacuum conditions on 18 flight units up to now has shown that the best focus will be between C and C, with an average value around C, well within the operational limits between C and C. The precise value for each camera will be determined during the commissioning phase once in space.
The optical design of each camera [221] consists of a six lens system with a central Calcium Fluoride lens [40] close to the stop, and with a first aspheric lens protected, for thermal and radiation hardness purposes, by a Suprasil window. All the optical materials have been the subject of specific radhard studies [94] and the whole design is a fine tuning balance between radiation, thermal and mass budgets [222]. To obtain more information on the current status of the point spread function analysis and the camera focusing activities, the interesting reader can refer to [48] and [265].
An anti-reflection coating is placed on every optical surface to maximise the optical transmission in the spectral range [500-1000] nm. Each 12 cm aperture camera has a FoV of 1037 square degrees as shown in [271]. This wide FoV is achieved thanks to a circular optical FoV of the TOU of around 18.88° radius that illuminates the 4 full-frame CCDs of 81.18 81.18 mm² each on the Focal Plane Assemblies (FPA). As mentioned above, each FPA is electrically connected to a Front End Electronics (either "normal" or "fast" ones), completing the PLATO Camera. One of the FPA models used for qualification is shown in Fig. 28, with 4 full-frame (i.e. "normal" CCDs). More details on the FPA can be found in Moreno et al [247]. Table 9 summarises the the main parameters of the individual cameras. The transmission and quantum efficiency curves can be consulted in [47].
@@ 13.3 PLATO Data Processing System
The PLATO payload on-board data processing system is made up of an Instrument Control Unit (ICU), two Main Electronic Units (MEU) and a Fast Electronic Unit (FEU). Each MEU contains 6 normal Data Processing Units (N-DPUs) for processing data from the normal cameras. Each N-DPU is dealing with two normal cameras in terms of managing the N-FEE and acquiring data. The nominal processing cadence for the N-DPUs is 25 sec. There are two F-DPUs gathered in the FEU. Each F-DPU is responsible for processing the data of one fast camera. The processing cadence for F-DPUs is 2.5 sec. The F-DPUs main purpose is providing attitude data for the Fine Guidance System (FGS) directly to the Service Module (SVM) AOCS and to the ICU. Furthermore, it manages the F-FEE and handles scientific data of bright stars. There are two ICU channels which work in cold redundancy. The ICU is responsible for the management of the payload, the communication with the SVM and the compression of scientific data before transmitting them as telemetry to the SVM. Data is routed through a SpaceWire network from FEU and MEU to the ICU, and then from the ICU to the SVM. For the FGS data, dedicated SpaceWire links between FEU and spacecraft AOCS are used. Figure 29 gives an overview of the PLATO data processing system architecture and of the data flow rates while Fig. 30 shows the engineering models during tests. It focuses on the sharing of the main functions and the data flows. It is a simplified view of the hardware architecture.
@@ 14 Synergies with other Missions
Exoplanet missions form a significant part of ESA and NASA’s ongoing and future space science programs ([309] for an overview). PLATO will start its science operations in 2027. By this time, the now ongoing extended operations periods for CHEOPS and TESS will have ended, but further prolongations could extend their lifetimes until and beyond PLATO launch. Prolongation of James-Webb-Space-Telescope (JWST) operations beyond 2026 will allow it to observe PLATO targets. Two years after PLATO, the ARIEL mission is set to launch and plans to spectroscopically observe transiting exoplanets. Here we outline the key aspects of synergies with PLATO for these and other missions:
@@
TESS: A discussion of the planet yield of PLATO versus TESS [282] can be found in Section 2. The majority of planets detected by TESS have orbital periods 10 days [147], while PLATO planets focus on orbital periods 27 days, including a significant fraction on long-periods 100 days. Most of the planet detections around solar-like stars from PLATO will have 4 REarth, outnumbering the numbers expected from TESS for this type of stars. There is a clear complementarity of TESS and PLATO. We note that if TESS enters a third extended mission period from 2026-2029, it will overlap with PLATO. This offers the opportunity for dedicated complementary observations where beneficial. This could include, for example, a pre-characterization of the PLATO field for short-period planets to prolonge TTV measurement baselines.
CHEOPS: The 30 cm-aperture space telescope is a follow-up mission carrying out high-precision photometry of known bright exoplanetary systems [39, 124]. CHEOPS has been used primarily to measure precise planetary radii, revising the composition of small planets [e.g, 45, 199, 114], to confirm transiting long-period planets [e.g., 106, 261], characterise multiplanet systems [e.g, 205, 220] and has also been used to characterise the emissive and reflective properties of planetary atmospheres [e.g, 206, 53].
@@
A major benefit to PLATO from CHEOPS results from improvements made in the planetary composition and structure modelling that inform the interpretation of precise planet radii and masses in terms of, e.g, planet formation, evolution and interior-atmosphere coupling. While many of the related science questions can only be fully answered when a larger sample of accurately characterised planets from PLATO becomes available, the ongoing activities within CHEOPS pave the way for PLATO data to come. With large parts of the PLATO LOPS2 field accessible to CHEOPS, a cross-calibration of CHEOPS and PLATO data of the same transits will benchmark the calibration of PLATO data (e.g. in terms of contamination removal). Preparatory observations of known transiting planets in the PLATO LOPS2 filed by CHEOPS have started and will be beneficial to maximising science return on these objects, crucially extending the baseline of transit observations for TTV studies.
Ariel: will provide spectra of the atmospheres of a large number (around 1000) of transiting planets [316]. These planets will mostly be hot and warm gaseous exoplanets (Jupiters, Saturns, Neptunes) as well as of super-Earths/sub-Neptunes orbiting bright stars of various types, with FG stars being the predominant host. Several surveys are already identifying potential Ariel targets that meet its mission objectives. For instance, TESS is expected to discover over 4500 planets around bright stars and more than 10000 giant planets around fainter stars (see Section 2). Other surveys, including PLATO, are expected to provide thousands of aditional candidates suitable for study with Ariel. Notably, about 20-40 targets are expected to come from a PLATO LOPs (see also Nascimbeni et al. 2024, submitted). Moreover, PLATO will likely uncover new small, warm planets around bright stars, providing an exciting pool of additional candidate targets for Ariel. PLATO will play a crucial role by delivering high-precision planetary radii, extended TTV datasets, and refined ephemerides for planets within its fields. Data from PLATO’s N-CAM (white-light phase curves) and F-CAMs (two-colour information) will further aid in prioritizing Ariel candidates, particularly for hot, nearby targets (see Section 3.3). The combined insights on radii, masses, and stellar parameters - especially stellar ages - will significantly enhance the scientific output of Ariel. For example, precise ages from PLATO will help break degeneracies in the study of warm Jupiters, enabling detailed investigations of planetary evolution, including contraction, atmospheric loss, and tidal interactions. In addition, PLATO’s pre-characterization of stellar activity (e.g., rotational modulation, activity cycles, flare and CME frequencies, granulation) will be invaluable. These data will be particularly useful for targets observed simultaneously by both missions.
@@
GAIA: Through its astrometric survey of the Milky Way, Gaia [128] will discover a large population of massive planets (expected to be super-Neptunes and more massive), many of which will be at long periods. Most of these will be found from their trigonometric signatures but some may also be detectable through their transits. A significant harvest of new long period ( 5 years), massive exoplanets is expected in the upcoming Gaia data releases (e.g. Gaia DR4). For those with sufficient orbital information, inclinations could be accurate enough to predict potential transit times and if these fall with the PLATO observation fields monitoring observations will be possible at low cost. These long period, transiting planets, would be extremely valuable. While many of these systems may only transit once during the PLATO observations, combined modelling of the astrometric and transit observations will produce accurate orbital and planetary parameters. For PLATO itself, the GAIA observations not only form the basis of the PLATO input catalogue, but allow contamination with the PLATO targets/pixels to be estimated. Furthermore, while GAIA produces limited temporal coverage it may reveal the presence of line of sight background eclipsing binaries as has been demonstrated for TESS candidates by [266]. This vetting capability will be extremely important in keeping the PLATO ground based followup tractable.
Euclid and Roman Space telescope: The Nancy Grace Roman Space telescope [308] is expected to detect exoplanets from microlensing, from transits and from coronographic imaging, covering a wide range of planet system parameters. ESA’s Euclid mission also has some capability to detect extrasolar planets, and studies demonstrating the joint detection capabilities of the two missions have been performed (e.g. Bachelet et al 20, Kerins et al 185). In particular the microlensing technique used by these missions has the potential to detect planets as small as Earth, and smaller, at orbital distances of 1 AU, although not with accurate planet parameters as anticipated by PLATO. Nevertheless, PLATO together with Roman and Euclid will be help solidify our understanding of the frequency of earth-sized planets in earth-like orbits. Both Euclid and the Roman Space Telescope will study a relatively distant stellar and hence planetary population which will contrast with the nearby PLATO planets. The large number of planets detected by these surveys will enable biases in their planet catch to be understood, hence leading to a better understanding of the planetary demographics in these different environments.
For all the missions discussed here, a combined analysis of their planetary frequencies will enable the most comprehensive picture of planetary demographics to be determined, allowing us to deduce how planets are born and evolve (through modelling).
@@ Acknowledgments
The authors gratefully thank David Ciardi and an unknown referee for carefully reading this publication and providing many helpful comments for clarifying the current status of PLATO to a large readership.
@@
M.L. acknowledges support of the Swiss National Science Foundation under grant number PCEFP2_194576. The contribution of M.L. has been carried out within the framework of the NCCR PlanetS supported by the Swiss National Science Foundation under grants 51NF40_182901 and 51NF40_205606.
@@
CdB acknowledges support from a Beatriz Galindo senior fellowship (BG22/00166) from the Spanish Ministry of Science, Innovation and Universities.
@@ 15 Appendix A
where describes the effect of limb darkening. The current knowledge of limb darkening for low magnetic field, quiet, solar-type stars is known sufficiently well that the contribution to the uncertainty in the radius ratio is negligible (c.f. Section 4.4. of Maxted 230 as well as Ludwig et al 217). However, numerical simulations show that about over 800 Gauss surface magnetic field the theory predicts higher limb darkening than that was observed. In the same way, some simulations result in limb brightening effects difficult to understand at this point (see Ludwig et al 217). So, the very active stars can have different properties and more work is needed to calibrate properly their limb darkening behaviour (c.f. Csizmadia et al 96, Espinoza and Jordán 119, Agol et al 7, Patel and Espinoza 269). The results are not changing if we consider different limb darkening laws; therefore, for sake of simplicity, we used linear limb darkening with a constant linear limb darkening coefficent which is a solar-like value (). Then we have by differentiation that